If AB = 2x+y, BC = 3x-2y, DC= x+2y+2 and AD= x+2y, what values of x and y will make ABCD a parallelogram?
Answer by Boreal(15235)

AB=CD so 2x+y=x+2y+2
x-y=2
BC=AD so 3x-2y=x+2y
2x=4y
so x=2y
but x-y=2
so y=2
x=4